# Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization

### v2

@@ Abstract
Rapid advancements in Large Language Models (LLMs) have accelerated their integration into automated visualization code generation applications. Despite advancements through few-shot prompting and query expansion, existing methods remain limited in handling ambiguous and complex queries, thereby requiring manual intervention. To overcome these limitations, we propose VisPath: a Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation. VisPath handles underspecified queries through structured, multi-stage processing. It begins by reformulating the user input via Chain-of-Thought (CoT) prompting, which refers to the initial query while generating multiple extended queries in parallel, enabling the LLM to capture diverse interpretations of the user intent. These queries then generate candidate visualization scripts, which are executed to produce diverse images. By assessing the visual quality and correctness of each output, VisPath generates targeted feedback that is aggregated to synthesize an optimal final result. Extensive experiments on widely-used benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath outperforms state-of-the-art methods, offering a more reliable solution for AI-driven visualization code generation.
Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
Wonduk Seo1,2   Seungyong Lee   Daye Kang1,3 Hyunjin An1   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author. 1Enhans   2Peking University   3KAIST   4Tsinghua University 1{wonduk,daye,hyunjin,seunghyun}@enhans.ai
@@ 1 Introduction
Figure 1: Overview of different approaches for visualization code generation. Comparing two baseline methods, namely Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al. (2024), with our proposed VisPath framework.
@@
Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets Vondrick et al. (2013); Demiralp et al. (2017); Unwin (2020); Li et al. (2024a). Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js Barrett et al. (2005); Bisong and Bisong (2019); Zhu (2013). These approaches demand programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al. (2018); Sharif et al. (2024). As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible Wang et al. (2015); Dibia and Demiralp (2019); Qian et al. (2021).
In response to this challenge, Large Language Models (LLMs) have emerged as a promising solution for simplifying visualization creation Wang et al. (2023a); Han et al. (2023); Xie et al. (2024). By translating natural language instructions into executable code, LLM-based systems eliminate the need for extensive programming knowledge, allowing users to generate visualizations more intuitively Xiao et al. (2023); Ge et al. (2023); Zhang et al. (2024b). Recent visualization methods such as ChartGPT Tian et al. (2024) and NL4DV Sah et al. (2024) demonstrate the potential of LLMs to provide interactive, conversational interfaces for visualization. These systems enable users to create complex charts with minimal effort, bridging the gap between technical expertise and effective data exploration Dibia (2023); Kim et al. (2024).
@@
More recently, LLM-based visualization frameworks such as Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al. (2024) have been introduced to improve automated visualization code generation. Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation. However, these methods face several limitations: (1) they generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; (2) they rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries; (3) they encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback. Without a mechanism to retrieve outputs that reflect diverse possibilities, they face difficulties in capturing the intricate details required for visualizations that are both functionally precise and contextually relevant.
To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a novel approach that redefines how visualization code is generated, as illustrated in Figure 1. Traditional methods often fall short of delivering the depth and precision users truly need, struggle to capture the intricate details that make a visualization not just functional but meaningful. VisPath challenges this limitation by incorporating Multi-Path Reasoning and Feedback-Driven Optimization, systematically exploring multiple interpretative pathways to construct a more accurate, context-aware, and fully executable visualization.
@@
Rather than simply translating user input into code, VisPath ensures that every critical aspect, both explicitly stated and implicitly necessary, is carefully considered, creating a visualization that is not only correct but insightful. At its core, it generates multiple reasoning paths that analyze the user’s intent from different perspectives, producing structured blueprints that are then transformed into visualization scripts through Chain-of-Thought (CoT) prompting. These multiple candidates are evaluated using a Vision-Language Model (VLM) to assess accuracy, clarity, and alignment with the intended message. The results are then refined through an Aggregation Module, optimizing the final output for both reliability and impact.
Extensive experiments on benchmark datasets, including MatPlotBench Yang et al. (2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate that VisPath notably outperforms state-of-the-art methods across all evaluation metrics. In addition, ablation studies confirm that VisPath’s performance gains primarily stem from its Multi-Path Reasoning mechanism and feedback-driven optimization. By systematically generating and evaluating multiple reasoning paths and leveraging feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents. Our investigations demonstrate its ability to capture nuanced user intents, improve execution reliability, and minimize errors, making visualization code generation more accessible and effective for domains such as business intelligence, scientific research, and automated reporting.
@@ 2 Related Work
Numerous methods have been applied for Text-to-Visualization (Text2Vis) generation, which has significantly evolved over the years, adapting to new paradigms in data visualization and natural language processing Dibia and Demiralp (2019); Wu et al. (2022); Chen et al. (2022a, b); Rashid et al. (2022); Zhang et al. (2024a). Early approaches such as Voyager Wongsuphasawat et al. (2015) and Eviza Setlur et al. (2016) largely relied on rule-based systems, which mapped textual commands to predefined chart templates or specifications through handcrafted heuristics de Araújo Lima and Diniz Junqueira Barbosa (2020). While these methods demonstrated the feasibility of automatically converting text into visualizations Moritz et al. (2018); Cui et al. (2019), they often required extensive domain knowledge and struggled with more nuanced or ambiguous user requirements Li et al. (2021); Wang et al. (2023b). Inspired by developments in deep learning, researchers began to incorporate neural networks to handle free-form natural language and broaden the range of supported visualization types Liu et al. (2021); Luo et al. (2021).
Figure 2: Overview of the proposed VisPath framework for creating robust visualization code generation. The framework consists of combination of Multi-Path Agent, Visual Feedback Agent, and Synthesis Agent.
Building on these machine learning strategies, numerous studies have utilized LLMs to further enhance system flexibility. Recent frameworks such as Chat2VIS Maddigan and Susnjak (2023) and Prompt4Vis Li et al. (2024b) utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting. More recent approaches, such as MatPlotAgent Yang et al. (2024) and PlotGen Goswami et al. (2025), extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations. The aforementioned approaches often struggle to effectively capture user intent in complex visualization tasks. By committing to a single reasoning trajectory, they may produce code that is syntactically correct yet semantically misaligned with user expectations, requiring extensive manual adjustments. This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification, consequently limiting the intended efficiency of automation. To address these limitations, we introduce VisPath, a novel framework that integrates Multi-Path Reasoning with feedback from VLMs to enhance visualization code generation.
@@ 3 Methodology
We introduce VisPath, a framework for robust visualization code generation that leverages diverse reasoning and visual feedback. VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a VLM evaluates and refines the outputs to ensure generation robust visualization code. An overview of this process is shown in Figure 2.
@@ 3.1 Multi-Path Generation
The potential for rigid interpretation is a major limitation in visualization code generation, a single query can have multiple valid representations depending on its dataset structure. VisPath mitigates this by generating multiple extended queries within a single interaction. Given a user query $Q$ and a corresponding dataset description $D$, a Multi-Path Agent is employed to expand the query into $K$ distinct reasoning pathways:
$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}=f_{\_}{\text{mpa}}(Q,D),$ (1)
where $f_{\_}{\text{mpa}}$ denotes the function of the Multi-Path Agent implemented via an LLM. The dataset description $D$ plays a crucial core in shaping these reasoning paths by providing contextual information about variable types, inherent relationships, and the suitability of different chart types for a more grounded interpretation of the query.
Each $R_{\_}i$ serves as a detailed logical blueprint outlining one possible approach to fulfill the visualization request. The purpose of generating multiple reasoning paths is not simply to increase their quantity, but to ensure interpretive diversity. Vispath mitigates the risk of depending on a single, potentially incorrect assumption by intentionally exploring alternative and context-aware reasoning strategies. This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent even when queries are ambiguous or underspecified.
3.2 Code Generation from Reasoning Paths
Once diverse reasoning paths are established, the subsequent stage involves translating each path into executable Python scripts. For each reasoning path $R_{\_}i$ generated in Equation 1, a dedicated Code Generation LLM produces the corresponding visualization code using Chain-of-Thought (CoT) prompting:
$C_{\_}i=f_{\_}{\text{code}}(D,R_{\_}i),$ (2)
where $f_{\_}{\text{code}}$ represents the code generation function. The dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes. The generated code $C_{\_}i$ is then executed to render a candidate visualization:
$V_{\_}i=f_{\_}{\text{exec}}(C_{\_}i),\quad i=1,2,\dots,K,$ (3)
with $f_{\_}{\text{exec}}$ serving as the code execution function. By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage, as the code has already been conditioned on $D$ during its generation.
In practice, some generated codes may not be executable. Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
$\epsilon_{\_}i=\begin{cases}1,&\text{if }C_{\_}i\text{ is executable}\\ 0,&\text{otherwise}.\end{cases}$ (4)
To route the outputs appropriately, we introduce:
$Z_{\_}i=\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1,\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0.\end{cases}$ (5)
The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the initial query $Q$, to the feedback model in the subsequent stage.
@@ 3.3 Feedback-Driven Code Optimization
While most code generation frameworks primarily focus on producing syntactically correct scripts, our framework extends beyond this by incorporating an additional mechanism. As final stage, VisPath synthesizes the most robust and accurate visualization code by leveraging both the executability information and structured visual feedback. A VLM is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$. This evaluation is formalized as:
$F_{\_}i=f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i),$ (6)
where $F_{\_}i$ provides structured feedback on key aspects such as chart layout, the alignment between the intended request and the rendered visualization (or error context), and visual readability (including potential improvements). To capture the complete quality signal from each candidate, we pair the feedback with its corresponding generated code:
$S_{\_}i=(C_{\_}i,F_{\_}i),\quad i=1,2,\dots,K.$ (7)
Leveraging the collective code-feedback pairs along with the initial query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
$C^{*}=f_{\_}{\text{integrate}}\left(Q,D,\{S_{\_}i\}_{\_}{i=1}^{K}\right),$ (8)
where $C^{*}$ represents the optimized visualization code and $f_{\_}{\text{integrate}}$ denotes the function that aggregates the strengths of each candidate code alongside its corresponding feedback. This formulation ensures that the final code is not only constructed based on the insights extracted from the candidate outputs but is also meticulously aligned with the original user query and the provided dataset description. An algorithm for VisPath is detailed in Algorithm 1 as shown below:
Algorithm 1 Algorithm for VisPath
1:User query $Q$, dataset $D$, number of reasoning paths $K$, Multi-Path Agent $f_{\_}{\text{mpa}}$, Code Generation LLM $f_{\_}{\text{code}}$, Code Execution Function $f_{\_}{\text{exec}}$, Feedback Model $f_{\_}{\text{feedback}}$, Integration Module $f_{\_}{\text{integrate}}$
2:// Step 1: Multi-Path Query Expansion
3:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$
4:// Step 2: Code Generation from Reasoning Paths
5:for $i=1$ to $K$ do
6:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$
7:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$
8:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$
9:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$
10:end for
11:// Step 3: Feedback-Driven Code Optimization
12:for $i=1$ to $K$ do
13:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$
14:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$
15:end for
16:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$
17:return $C^{*}$
@@ 4.1.1 Experimental Datasets
We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al. (2024) and Qwen-Agent Code Interpreter Benchmark. Specifically, MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries. In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.

### v3

@@ Abstract
Large Language Models (LLMs) have become a cornerstone for automated visualization code generation, enabling users to create charts through natural language instructions. Despite improvements from techniques like few-shot prompting and query expansion, existing methods often struggle when requests are underspecified in actionable details (e.g., data preprocessing assumptions, solver or library choices, etc.), frequently necessitating manual intervention. To overcome these limitations, we propose VisPath: a Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation. VisPath handles underspecified queries through structured, multi-stage processing. It begins by using Chain-of-Thought (CoT) prompting to reformulate the initial user input, generating multiple extended queries in parallel to surface alternative plausible concretizations of the request. These queries then generate candidate visualization scripts, which are executed to produce diverse images. By assessing the visual quality and correctness of each output, VisPath generates targeted feedback that is aggregated to synthesize an optimal final result. Extensive experiments on MatPlotBench and Qwen-Agent Code Interpreter Benchmark show that VisPath outperforms state-of-the-art methods, providing a more reliable framework for AI-driven visualization generation.
@@ 1 Introduction
Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets [28, 9, 27, 14]. Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js [2, 3, 40]. These approaches demand programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users [4, 23, 25]. As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible [31, 10, 21].
@@
Large Language Models (LLMs) have emerged as a promising solution for simplifying visualization creation [29, 13, 36, 32]. By translating natural language instructions into executable code, LLM-based systems eliminate the need for extensive programming knowledge, allowing users to generate visualizations more intuitively [35, 11, 39]. More recently, Chat2VIS [19] and MatPlotAgent [37] have been introduced to improve automated visualization code generation. Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation. However, these methods face several limitations: (1) they generate code in a single-path manner, limiting exploration of alternative solutions, and are unable to recover when generating erroneous code; (2) they rely on predefined structures or examples, which restrict adaptability to ambiguous or unconventional user queries; and (3) a fundamental limitation is their inability to aggregate and synthesize multi-dimensional feedback. Without a mechanism to retrieve outputs that reflect diverse possibilities, they face difficulties in capturing the intricate details required for visualizations that are both functionally precise and contextually relevant.
To address underspecified visualization requests, we propose VisPath, a Multi-Path Reasoning and Feedback-Driven Optimization framework. VisPath samples multiple plausible concretizations of implicit implementation choices, executes candidate scripts, and leverages visual feedback to synthesize a robust final program.
@@
Rather than directly translating user input into code, VisPath systematically accounts for both explicit requirements and implicit necessities to produce visualizations that are correct and insightful. It generates multiple reasoning paths that interpret user intent from different perspectives, producing structured blueprints that are converted into visualization scripts via Chain-of-Thought prompting. These candidates are then evaluated by a Vision-Language Model (VLM) for accuracy, clarity, and alignment with the intended message, and refined by a synthesis agent to optimize reliability and impact.
Experiments show that VisPath improves plot-level correctness and executability over prompting and agent-based baselines. Ablations attribute the gains to multi-path exploration and feedback integration, demonstrating stronger intent capture, higher execution reliability, and fewer errors—making visualization code generation more accessible for business intelligence, scientific research, and automated reporting.
@@ 2 Related Work
Numerous methods have been applied for Text-to-Visualization (Text2Vis) generation, which has significantly evolved over the years, adapting to new paradigms in data visualization and natural language processing [10, 34, 5, 6, 22, 38]. Early approaches such as Voyager [33] and Eviza [24] largely relied on rule-based systems, which mapped textual commands to predefined chart templates or specifications through handcrafted heuristics [8]. While these methods demonstrated the feasibility of automatically converting text into visualizations [20, 7], they often required extensive domain knowledge and struggled with more nuanced or ambiguous user requirements [15, 30]. Inspired by developments in deep learning, researchers began to incorporate neural networks to handle free-form natural language and broaden the range of supported visualization types [17, 18].
Building on these machine learning strategies, numerous studies have utilized LLMs to further enhance system flexibility. Recent frameworks such as Chat2VIS [19] and Prompt4Vis [16] utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting. More recent approaches, such as MatPlotAgent [37] and PlotGen [12], extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations. The aforementioned approaches often struggle to effectively capture user intent in complex visualization tasks. By committing to a single reasoning trajectory, they may produce code that is syntactically correct yet semantically misaligned with user expectations, requiring extensive manual adjustments. This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification, consequently limiting the intended efficiency of automation. To address these limitations, we introduce VisPath, a novel framework that integrates Multi-Path Reasoning with feedback from VLMs to enhance visualization code generation.
@@ 3 Methodology
We introduce VisPath, a framework for robust visualization code generation that combines diverse reasoning with visual feedback. Given a user query $Q$ and a dataset description $D$, VisPath proceeds in three stages: (i) multi-path query expansion, (ii) visualization code synthesis and execution, and (iii) feedback-driven refinement. An overview is shown in Figure 1.
Figure 1: Overview of the proposed VisPath framework for creating robust visualization code generation. The framework consists of a combination of Multi-Path Agent, Visual Feedback Agent, and Synthesis Agent.
Notation.
We denote any agent component by $G_{\bullet}(\,\cdot\mid S_{\bullet})$, where $S_{\bullet}$ is the system prompt that specifies the role, constraints, and output format for that component. In contrast, code execution is performed by a non-Agent operator, denoted $\mathcal{E}(\cdot)$.
@@ 3.1 Multi-Path Generation
A single visualization request can support multiple valid interpretations depending on dataset schema, variable types, and implicit relationships. To mitigate brittle, single-assumption interpretations, VisPath generates a diverse set of reasoning paths conditioned on the dataset context. Specifically, a Multi-Path Agent produces $K$ distinct reasoning pathways:
$\{R_{1},R_{2},\dots,R_{K}\}\sim G_{\text{mpa}}(Q,D\mid S_{\text{mpa}}).$ (1)
Here each $R_{i}$ is a structured plan (a logical blueprint) describing one plausible data-grounded interpretation of $Q$ under $D$ (e.g., chart type choice, grouping/aggregation assumptions, encoding decisions, and potential preprocessing). Conditioning on $D$ encourages paths that respect variable names, data types, and feasible visual encodings.
3.2 Code Generation and Execution
For each reasoning path $R_{i}$, a code-generation agent synthesizes an executable Python visualization script:
$C_{i}\sim G_{\text{code}}(D,R_{i}\mid S_{\text{code}}),\quad i=1,\dots,K.$ (2)
The dataset description $D$ is provided explicitly to ground the script in the actual data context (e.g., column names, semantic types, and plausible transformations), reducing hallucinated variables and mismatched encodings.
Each generated script is then executed by a deterministic execution operator $\mathcal{E}$. We define the routed execution output as a tuple that jointly encodes executability and the observable outcome:
$Z_{i}:=\mathcal{E}(C_{i})=\begin{cases}\bigl(1,\,\mathrm{plot}(V_{i})\bigr),&\text{if }C_{i}\text{ executes successfully and renders }V_{i},\\ \bigl(0,\,\mathrm{err}(C_{i})\bigr),&\text{otherwise},\end{cases}$ (3)
for $i=1,\dots,K$, where the first element indicates executability and the second element is either the rendered plot image or the execution error message. This design avoids an explicit debugging loop while still preserving informative failure signals for downstream refinement.
@@ 3.3 Feedback-Driven Code Optimization
Next, a feedback agent evaluates each candidate by jointly considering the user intent and the observed execution outcome:
$F_{i}\sim G_{\text{fb}}(Q,C_{i},Z_{i}\mid S_{\text{fb}}),\quad i=1,\dots,K,$ (4)
where $F_{i}$ is structured feedback capturing (i) semantic alignment with $Q$, (ii) correctness with respect to the data context implied by $D$, and (iii) visual quality/readability (e.g., labeling, legends, scales, occlusion, layout).
Finally, a Synthesis Agent synthesizes a refined visualization program by aggregating the full set of candidates and their feedback:
$C^{*}\sim G_{\text{syn}}\!\left(Q,D,\{(C_{i},F_{i})\}_{i=1}^{K}\mid S_{\text{syn}}\right).$ (5)
The output $C^{*}$ is optimized to be (1) executable, (2) faithful to the user intent in $Q$, and (3) visually informative given the dataset context $D$, by selectively inheriting strengths and correcting weaknesses identified across the candidate set.
Summary.
VisPath generates multiple dataset-aware reasoning paths, translates them into candidate scripts, executes each script to obtain either a plot or an error signal, and then uses structured multimodal feedback to synthesize these signals into a final robust visualization program.
@@ 4.1.1 Experimental Datasets
We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench [37] and Qwen-Agent Code Interpreter Benchmark. Specifically, MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries. In this paper, we use underspecification to mean that a query omits at least one decision that must be instantiated in code (e.g., plot construction primitives, data transformations, or layout conventions). Although each benchmark item provides a single reference visualization, many prompts still require such implicit choices. For example, a broken-axis plot typically entails a two-panel shared-x layout with different y-limits, and polar bar charts require a polar projection with angles in radians. VisPath is designed to surface and resolve these implicit choices via multi-path concretization. The Qwen-Agent Code Interpreter Benchmark contains 295 records, of which 163 are visualization-related, and evaluates Python code-interpreter agents using Code Executability and Code Correctness on tasks including data visualization.